OpenStudy (anonymous):

A cone has a base radius of 2 millimeters and a height of 24 millimeters. What is the volume of the cone? A. 16π mm3 B. 96π mm3 C. 384π mm3 D. 32π mm3

OpenStudy (phi):

do you know the formula for the volume of a cone ?

OpenStudy (anonymous):

yea v=1/3TTr^2h

OpenStudy (phi):

ok, \[ V = \frac{1}{3} \pi r^2 h \] replace h with 24 replace r with 2 can you do that ?

OpenStudy (anonymous):

yeah

OpenStudy (phi):

can you post it ?

OpenStudy (anonymous):

V=1/3TT(2)^2(24)

OpenStudy (phi):

remember order of operations. we do exponents first what is 2^2 ?

OpenStudy (anonymous):

4

OpenStudy (phi):

so replace 2^2 with 4 V=1/3TT(4)(24) that says multiply 1/3 * pi * 4 * 24 we can multiply in any order. put the pi to the aside, and do the numbers

OpenStudy (phi):

I would first do 1/3 * 24 (because you get a nice number)

OpenStudy (anonymous):

V=1/3TT*4*24 v=1/3*96 V=96/3TT

OpenStudy (anonymous):

IS that right

OpenStudy (phi):

if you mean \[ V= \frac{96}{3} \pi\] yes. but 3 divides evenly into 96 (I would have done 24/3 = 8 first, and then 8*4= 32) but you can do it this way what is 96/3 ?

OpenStudy (anonymous):

okay thankyou

OpenStudy (phi):

what did you get ?

OpenStudy (anonymous):

32

OpenStudy (phi):

yes, so \[ V= 32 \pi\] btw, the reason they did not multiply pi * 32 is because pi is 3.14159265358979323846…(goes on forever) and rather than pick some "rounded version" such as 3.14 they leave it 32 pi then if you want a number, you use pi to the accuracy you want.

OpenStudy (anonymous):

okay thanks for the help